3/6 = 5x can you show me the operation on gettin answer? Tks
Found 2 solutions by linh117, Adrish:
Answer by linh117(31)

(3)/(6)=5x
Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
5x=(3)/(6)
Reduce the expression (3)/(6) by removing a factor of 3 from the numerator and denominator.
5x=(1)/(2)
Divide each term in the equation by 5.
(5x)/(5)=(1)/(2)*(1)/(5)
Simplify the left-hand side of the equation by canceling the common terms.
x=(1)/(2)*(1)/(5)
Simplify the right-hand side of the equation by simplifying each term.
x=(1)/(10)

Answer by Adrish(44)

Solutions:
3/6 = 5x
Write the fraction in the reduced and simplest form
1/2=5x
Multiply both sides by 1/5, we get
1/2 x 1/5= 5x (1/5)
x=1/10